# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ Abstract
Preference learning (PL) with large language models (LLMs) aims to align the LLMs’ generations with human preferences. Previous work on reinforcement learning from human feedback (RLHF) has demonstrated promising results in in-distribution PL. However, due to the difficulty of obtaining human feedback, discretely training reward models for every encountered distribution is challenging. Thus, out-of-distribution (OOD) PL is practically useful for enhancing the generalization ability of LLMs with limited preference feedback. This work addresses OOD PL by optimizing a general reward model through a meta-learning approach. During meta-training, a bilevel optimization algorithm is utilized to learn a reward model capable of guiding policy learning to align with human preferences across various distributions. When encountering a test distribution, the meta-test procedure conducts regularized policy optimization using the learned reward model for PL. We theoretically demonstrate the convergence rate of the bilevel optimization algorithm under reasonable assumptions. Additionally, we conduct experiments on two text generation tasks across 20 held-out domains and outperform a variety of strong baselines across various evaluation metrics.
@@ 1 Introduction
Aligning large language models (LLMs) with human preferences through reinforcement learning (RLHF) has been demonstrated as a practical approach to align pretrained LLMs along human values. As demonstrated by recent research on LLMs (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022), RLHF initially trains a reward model to capture human preferences from a dataset of pairwise comparisons. It then aligns the LLMs with the learned reward model through regularized policy optimization, aiming to learn a language policy that better reflects human values. However, defining precise rewards for various real-world tasks is non-trivial (McKinney et al., 2022), and obtaining high-quality feedback that accurately represents human preferences is challenging Bai et al. (2022).
@@
Most previous work on RLHF focuses on in-distribution preference learning (Casper et al., 2023). For example, when conducting sentiment generation for the film distribution, RLHF utilizes only the in-distribution preference data for reward learning and performs regularized policy optimization with the in-distribution data. However, two major challenges arise when encountering the distribution shift problem. For example, Figure 1 illustrates a scenario where preference data from the book and dvd domains are used for reward modeling, and regularized policy optimization is conducted with data from the book distribution. The first challenge is that the generations from the policy may substantially deviate from the human preferences of the target distribution, posing an out-of-distribution (OOD) challenge for preference learning. Furthermore, these distribution shifts result in policy drift during the policy optimization process with Kullback-Leibler divergence (Ramamurthy et al., 2023; Rafailov et al., 2024): the policy tends to move towards the reference policy largely followed in the training distribution, exacerbating the model’s deviation from the target distribution in preference learning.
@@
An initial approach to addressing the out-of-distribution (OOD) challenge is to directly enhance the OOD generalization ability of the reward model using preference data from multiple distributions. This specific approach may involve employing multi-task learning to train a reward model using data from various distributions or integrating reward models from multiple domains into a unified model, similar to previous work on reward model ensembling (Ramé et al., 2024). However, solely improving reward modeling for OOD reward prediction within the scope of preference learning may be suboptimal, considering that the objective of preference learning is to optimize a policy aligned with human preferences. To better tackle the OOD challenge, we employ an end-to-end strategy to learn a reward function capable of guiding policy optimization for OOD preference learning. To achieve this, we utilize a meta-learning approach to directly compute OOD preferences based on the predicted probabilities generated by a language policy. This policy is optimized w.r.t. a reward-model-based objective for policy optimization using in-distribution fine-tuning data. In particular, we incorporate a regularization term in policy optimization to penalize the Kullback-Leibler divergence between the predicted distribution by the policy and the sampling distribution of examples from the target distribution, thereby mitigating the challenge of policy drift.
To implement the above approach, we propose a gradient-based bilevel optimization algorithm. Specifically, the outer loop performs a sufficient number of stochastic gradient descent (SGD) iterations to achieve convergence and optimize the out-of-distribution (OOD) preference learning objective with respect to the reward model. The inner loop performs SGD updates for the in-distribution policy using the reward model. As illustrated in Figure 1, the entire training process comprises a meta-training phase and a meta-test phase. In the meta-test phase, a sufficient number of SGD iterations are conducted to optimize a language policy with the meta-learned reward model using supervised samples from the test distribution. Theoretically, we establish an upper bound for the convergence rate of bilevel optimization with respect to the learning rates of inner-loop and outer-loop optimization, as well as the controlling factor for reward modeling. Empirically, we conduct experiments on both controlled sentiment generation and knowledge answer generation, using multiple metrics for evaluation. The results demonstrate that our method outperforms a range of strong baselines for OOD preference learning and achieves the best results across four domains of sentiment generation and 18 domains of answer generation.
Figure 1: The meta-learning phase and testing phase are crucial for out-of-distribution preference learning. During the meta-training phase, the reward model $\Phi$ is optimized using $K$ iterations of stochastic gradient descent (SGD) on preference examples and fine-tuning samples from training distributions, such as the book and dvd distributions. Upon encountering an unseen distribution film during the testing phase, the optimized reward model $\Phi_{K}$ from the meta-training phase is utilized, and only the in-distribution fine-tuning samples are used to optimize the in-distribution policy $\theta_{*}$.
@@ 2 Preliminaries
Problem Setting Training Inference
preference data fine-tuning data
SFT - $\mathcal{D}^{tr}$, $\mathcal{D}^{te}$ $\mathcal{D}^{te}$
in-distribution PL $\mathcal{D}_{\rm PL}^{tr}$, $\mathcal{D}^{te}_{\rm PL}$ $\mathcal{D}^{tr}$, $\mathcal{D}^{te}$ $\mathcal{D}^{te}$
out-of-distribution PL $\mathcal{D}_{\rm PL}^{tr}$ $\mathcal{D}^{tr}$, $\mathcal{D}^{te}$ $\mathcal{D}^{te}$
Table 1: Problem Setting. $\mathcal{D}{\rm PL}^{tr}$ and $\mathcal{D}{\rm PL}^{te}$ represent preference data from the training and testing distributions, respectively, while $\mathcal{D}^{tr}$ and $\mathcal{D}^{te}$ denote fine-tuning data from the training and testing distributions, respectively.
Notations.
Building on the work of Azar et al. (2023); Rafailov et al. (2024), we define the preference learning objective as learning a language model policy $\pi_{\theta}$ parameterized by $\theta$. Specifically, given a language prompt $x$ drawn from the input distribution $\mathcal{D}_{X}$ over the input space $\mathcal{X}$, the policy computes the probability distribution $\pi(\cdot|x)$ over the answer space $\mathcal{Y}_{x}$ and the predicted answer is denoted as $\hat{y}=\mathop{\arg\max}_{y\in\mathcal{Y}_{x}}\pi_{\theta}(y|x)$. Traditional supervised fine-tuning uses prompt-response pairs $(x,y)$ drawn from a joint distribution $\mathcal{D}_{XY}$ for alignment learning. In this paper, we use $\mathcal{D}$ to represent both $\mathcal{D}_{X}$ and $\mathcal{D}_{XY}$ for brevity. To align with the human preferences for answer generation using pretrained language models, preference learning utilizes pairs of answers corresponding to a prompt, $\{y,y^{\prime},x:y\succ y^{\prime}|x\}$, drawn from a distribution $\mathcal{D}_{\rm PL}$ where $y\succ y^{\prime}|x$ indicates that $y$ is the preferred answer over dispreferred answer $y^{\prime}$ given the prompt $x$.
Supervised fine-tuning (SFT).
SFT involves fine-tuning a pretrained LLM to adapt to task-specific datasets. The SFT process involves employing a supervised learning objective, specifically maximum likelihood estimation (MLE), on a carefully curated high-quality dataset, denoted as $\mathcal{D}$. The training objective is to obtain a language policy $\pi_{\theta}$ that satisfies
$\min_{\theta}-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\log\pi_{\theta}(y|x)\right]$ (1)
RLHF with reward modeling.
The standard RLHF framework (Ziegler et al., 2019; Ouyang et al., 2022) for aligning human preferences with language models comprises two main phases: (i) learning a reward model for a specific task, and (ii) fine-tuning the policy with the learned reward model. During the reward model learning phase, given a pairwise data distribution $\mathcal{D}_{\rm PL}$, a reward model $r_{\phi}:(x,y)\mapsto r_{\phi}(x,y)\in\mathbb{R}$ parameterized by $\phi$ is trained to score preference of $y$ to be the answer of $x$, using a maximum likelihood estimation (MLE) loss function:
@@
where $\sigma$ denotes the logistic function. In the RL fine-tuning phase, given a prompt distribution $\mathcal{D}$, the optimization objective can be represented as follows,
$\max_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi_{\theta}(\cdot|x)||\pi_{\rm ref}(\cdot|x)\right)\right],$ (3)
where $\mathrm{D}{\rm KL}(\cdot||\cdot)$ represents the Kullback-Leibler divergence, $\pi{\rm ref}(\cdot|x)$ denotes a reference distribution, and $\beta$ is the balancing coefficient. Previous work, such as RLHF-PPO(Ramamurthy et al., 2023) and DPO Rafailov et al. (2024), uses an SFT model as the $\pi_{\rm ref}$ to impose constraints on the similarity between the predictions of the RLHF model and those of the initial SFT model. In this study, we revise the constraint using $\pi_{\rm ref}$ to accommodate meta-learning for generalization across domains, as detailed in §3.
Problem Setting.
We compare three problem settings for language policy learning, as shown in Table 1. To ensure a fair comparison, we consider a transfer learning scenario where all learning settings utilize supervised fine-tuning data from both the training and test distributions $\mathcal{D}^{tr}$, $\mathcal{D}^{te}$ during training, and employ prompting from the test distribution $\mathcal{D}^{te}$ during inference. In contrast to supervised fine-tuning (SFT), preference learning (PL) utilizes additional human preference data for training, as demonstrated in previous works (Christiano et al., 2017; Ouyang et al., 2022; Zhang et al., 2023; Liu et al., 2023; Rafailov et al., 2024). In the in-distribution PL setting, similar to previous approaches such as RLHF-PPO (Christiano et al., 2017) and DPO Rafailov et al. (2024), preference data from both the training and test distributions are used for training. This work focuses on an out-of-distribution PL setting where only preference data from the training distribution are used for training, and fine-tuning data from both the training and testing distributions are used for policy optimization.
@@ 3 Meta-Learning for Generalizing Reward Modeling
Bilevel Programming
Reward Optimization (in-distribution)
Meta-Learning (out-of-distribution)
outer variables
reward model $\phi$
generalized reward model $\phi$
outer objective
PL objective $\mathcal{L}_{\rm PL}(\phi,\theta_{*},\mathcal{D}_{\rm PL})$ (Eq. (7))
PL objective ${\bf L}(\phi)$ (Eq. (11))
inner variables
policy $\theta$
policies of each distribution $\{\theta^{T}\}_{T\in\mathcal{D}_{\mathcal{T}}}$
inner objective
FT objective $\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D})$ (Eq. (10))
FT objectives for each distribution $\{\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D}^{T})\}_{T\in\mathcal{D}_{\mathcal{T}}}$ (Eq. (12))
Table 2: Linkes and naming conventions among different fields of bilevel optimization in this work.
To develop a meta-learning approach for generalizing reward modeling, we employ bilevel programming, building on previous research on meta-learning for generalization (Finn et al., 2017; Franceschi et al., 2018; Li et al., 2018; Ji et al., 2021). In this section, we first introduce a bilevel optimization framework for optimizing rewards in §3.1. Following the bilevel optimization framework, we then introduce the meta-learning approach for OOD PL in §3.2. Table 2 outlines the connections among bilevel programming, in-distribution reward optimization, and meta-learning for OOD PL. Additionally, we introduce a gradient-based algorithm for optimizing the meta-learning objective in §3.3 and analyze the theoretical convergence rate in §3.4.
3.1 Reward Optimization via Bilevel Programming
Bilevel programming.
In this section, we consider a bilevel optimization problem (see e.g., (Colson et al., 2007; Sinha et al., 2017)) of the form:
@@ 3.1 Reward Optimization via Bilevel Programming
${\rm s.t.}\ \theta_{*}\in\mathop{\arg\min}_{\theta}\mathcal{L}_{\rm FT}({\phi},{\theta};\mathcal{D}),$ (5)
@@
where $\mathcal{L}_{\rm PL}$ represents the outer objective, and for every reward function $r_{\phi}$, $\mathcal{L}_{\rm FT}$ represents the inner objective. Note that $\theta_{*}$ denotes the optimal parameters of a policy w.r.t. a reward model $r_{\phi}$ by RLHF fine-tuning via minimizing the inner objective $\mathcal{L}_{\rm FT}$.
Outer objective.
@@
In the context of reward optimization, we desire that the fine-tuned policy ${\theta}_{*}$, with the help of the reward function $r_{\phi}$, can align with human preferences. In particular, given a pair of answers for a prompt $x$ with preference ranking, $(y,y^{\prime},x:y\succ y^{\prime}|x)\sim\mathcal{D}_{\rm PL}$, we use the Bradley-Terry (BT) model (Bradley & Terry, 1952) to compute the comparing preference of $p(y\succ y^{\prime}|x)$ w.r.t. a fine-tuned policy ${\theta_{*}}$ as follows
$p_{*}(y\succ y^{\prime}|x)=\frac{\pi_{\theta_{*}}(y|x)}{\pi_{\theta_{*}}(y|x)+\pi_{\theta_{*}}(y^{\prime}|x)}$ (6)
@@
Given a preference sampling distribution (a.k.a. a dataset) $\mathcal{D}_{\rm PL}=\{z=(x,y,y^{\prime})\}$, we apply the maximum likelihood estimation (MLE) to define the outer objective of preference alignment w.r.t. the policy ${\theta_{*}}$ as follows
@@
$\begin{split}&\min_{\phi}\mathcal{L}_{\rm PL}(\phi,\theta_{*};\mathcal{D}_{\rm PL})=\min_{{\phi}}\mathbb{E}_{z\sim\mathcal{D}_{\rm PL}}\left[\ell_{\rm PL}({\phi},{\theta}_{*},z)\right]\\ &{\rm s.t.}\ \ell_{\rm PL}\left({\phi},{\theta}_{*},z\right)=-\log\sigma\left(\log\pi_{\theta_{*}}(y|x)-\log\pi_{\theta_{*}}(y^{\prime}|x)\right)\end{split}$ (7)
Inner objective.
@@
For the inner-loop optimization, we consider the objective of RLHF fine-tuning in Eq. (3), and specify the reference policy as the sampling distribution for the target task to avoid the policy drift problem that the learned policy may move towards the source distribution. In particular, given a target distribution $\mathcal{D}$, the reference distribution is represented as the sampling distribution of $\mathcal{D}$, i.e., $\pi_{\rm ref}(\cdot|x):=\hat{\mu}(\cdot|x)$ for any $x\sim\mathcal{D}$. Then, the RLHF fine-tuning objective can be represented as
@@
$\mathop{\arg\max}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi_{\theta}(\cdot|x)||\hat{\mu}(\cdot|x)\right)\right]$ (8)
@@
However, it is unsolvable to directly optimize the above objective using the policy gradient optimization approach Schulman et al. (2015; 2017), since the sampling distribution is unknown. Thus, we consider an approximate solution to this optimization objective. This derivation largely follows Peters & Schaal (2007); Peng et al. (2019). The major observation is that the objective is strictly concave and thus we can use the Karush-Kuhn-Tucker conditions (Boyd & Vandenberghe, 2004) to obtain a globally optimal solution to Eq. (8) such that for any $x\sim\mathcal{D}$ and $y\in\mathcal{Y}_{x}$,
$\pi_{*}(y|x)=\frac{1}{Z(x)}\hat{\mu}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right),$ (9)
@@
where $Z(x)$ represents the partition function. A detailed derivation is available in Appendix A. We focus on a parameterized estimator (e.g., a neural network) to tackle a parameter estimation problem $\min_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}}[\mathrm{D}_{\rm KL}(\pi_{*}(\cdot|x)||\pi_{\theta}(\cdot|x))]$. Thereby, given a fine-tuning sampling distribution $\mathcal{D}=\{z=(x,y)\}$, the inner objective can be represented as a logical regression formulaion,
@@
$\begin{split}&\mathop{\arg\min}_{\theta}\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D})=\mathop{\arg\min}_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\hat{\mu}(y|x)}\left[\ell_{\rm FT}({\phi},{\theta},z)\right]\\ &\qquad\qquad\qquad\qquad\quad=\mathop{\arg\min}_{\theta}\mathbb{E}_{z\sim\mathcal{D}}\left[\ell_{\rm FT}({\phi},{\theta},z)\right]\\ &{\rm s.t.}\ \ell_{\rm FT}({\phi},{\theta},z)=-\log\pi_{\theta}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\end{split}$ (10)
3.2 Meta-Learning

### v2

@@ Abstract
Preference learning (PL) with large language models (LLMs) aims to align the LLMs’ generations with human preferences. Previous work on reinforcement learning from human feedback (RLHF) has demonstrated promising results in in-distribution PL. However, due to the difficulty of obtaining human feedback, discretely training reward models (RMs) for every encountered distribution is challenging. Thus, out-of-distribution (OOD) PL is useful for enhancing LLMs’ generalization ability with limited preference feedback. This work addresses OOD PL by optimizing a general RM through a meta-learning approach. A bilevel optimization algorithm is utilized during meta-training to learn an RM that guides policy learning to align with human preferences across various distributions. When encountering a test distribution, the meta-test procedure optimizes regularized policy using the learned RM for PL. We theoretically demonstrate the convergence rate of the bilevel optimization algorithm under reasonable assumptions. Additionally, we conduct experiments on two text generation tasks across 22 held-out data distributions and outperform various strong baselines across various evaluation metrics.
Keywords:
Preference optimization LLMs OOD generalization.
@@ 1 Introduction
Aligning large language models (LLMs) with human preferences through reinforcement learning has been demonstrated as a practical approach to align pretrained LLMs along human values. As shown by recent research on LLMs [10, 37, 35, 26], RLHF initially trains a reward model (RM) to capture human preferences from a pairwise preference dataset. It then aligns the LLMs with the learned RM through regularized policy optimization, aiming to learn a language policy that better reflects human values. Most previous work on RLHF focuses on in-distribution (ID) preference learning (PL) [9], i.e., using the ID preference data for reward learning [10] and then performing policy optimization using PPO (Fig. 1 (b)). Besides, direct preference optimization (DPO) [30] and its variants [2, 15, 27] directly optimize a language policy using maximum likelihood estimation on the ID preference data (Fig. 1 (c)).
@@
However, defining precise rewards for various real-world tasks is non-trivial [24], and obtaining high-quality feedback that accurately represents human preferences is challenging [4]. Therefore, we focus on enhancing the out-of-distribution (OOD) generalization ability for PL. Fig. 1 (a) illustrates an OOD scenario where only preference data from training distributions ($y$,$y^{\prime}$,…, $y$,$y^{\prime}$) are available, as well as fine-tuning data from training-/test-distributions ($\mathcal{D}^{{tr}_{1}}_{xy}$,…,$\mathcal{D}^{{tr}_{N}}_{xy}$, $\mathcal{D}^{{te}}_{xy}$). Two major challenges arise when these ID PL methods encounter the distribution shift: (i) the generations from the policy may substantially deviate from the human preferences of the target distribution, posing an OOD challenge for PL. (ii) the distribution shifts result in policy drift during the policy optimization process with Kullback-Leibler divergence [31, 30]: the policy tends to move towards the reference policy largely followed in the training distribution, exacerbating the model’s deviation from the test distribution in PL.
Figure 1: Comparison between OOD PL and existing ID PL, i.e., PPO and DPO.
@@
An initial approach to addressing the OOD challenge is to directly enhance the OOD generalization ability of the RM using preference data from multiple distributions. This specific approach may involve training an RM using multiple distributions or integrating distribution-specific RMs into a unified model, similar to previous work on RM ensembling [32]. However, such a method also has limitations from the distribution shift between the preference data and fine-tuning data. To better tackle the OOD challenge, we employ a meta-learning approach to learn a reward function capable of guiding policy optimization for OOD preference learning (Fig. 1 (a)). For each episode, the policy is optimized with RL on fine-tuning data (meta-training on support set) and the RM is optimized with PL objectives on the preference data (meta-test on query set). In particular, we incorporate a regularization term in policy optimization to penalize the Kullback-Leibler divergence between the policy-induced distribution and test distribution, thereby mitigating the challenge of policy drift. We summarize the contributions as follows.
  • Improving OOD generalization for preference learning. We focus on aligning large language models with human preference on the OOD data, and propose a novel meta-learning-based approach to tackle the OOD preference learning problem.
  • Gradient-based bilevel optimization algorithm. To optimize the bilevel optimization objectives of meta-learning, we propose a gradient-based algorithm, and we establish an upper bound for its convergence rate w.r.t. the learning rates and controlling factor for reward modeling.
  • State-of-the-art performance. We conduct experiments on both controlled sentiment generation and knowledge answer generation, using multiple metrics for evaluation. The results outperform a range of strong baselines and achieve the best results across four sentiment generation distributions and 18 answer generation distributions. The code is released at https://github.com/jiachenwestlake/OODPL.
@@ 2 Preliminaries
Notations. We relay on an off-policy sequence-to-sequence decision making framework of language generation [3, 2, 30]: given a task-specific prompt $x\sim\mathcal{D}_{x}$, the $\theta$-parameterized language policy $\pi_{\theta}$ computes the probability distribution of response $y\sim\pi_{\theta}(\cdot|x)$, together with a task-specific sequence-level reward $r_{\phi}(x,y)\in\mathbb{R}$ by the $\phi$-parameterized RM $r_{\phi}$. Preference learning (PL) [10, 26, 30] further utilizes answer comparisons corresponding to a prompt, $\{y,y^{\prime},x:y\succ y^{\prime}|x\}$, drawn from a preference distribution $\mathcal{D}_{\rm PL}$ where $y\succ y^{\prime}|x$ indicates that $y$ is the preferred answer over dispreferred answer $y^{\prime}$ given the prompt $x$.
Table 1: Comparison of training data distributions. $\mathcal{D}_{\rm PL}^{tr}$ and $\mathcal{D}_{\rm PL}^{te}$ represent training and testing preference distributions, respectively, $\mathcal{D}^{tr_{i}}_{xy}$ denote the RL (or SFT) data distribution.
Method PL RL (SFT)
SFT (ID) - $\mathcal{D}^{te}_{xy}$
PPO (ID) $\mathcal{D}^{te}_{\rm PL}$ $\mathcal{D}^{te}_{xy}$
xDPO (ID) $\mathcal{D}^{te}_{\rm PL}$
ours (OOD) $\{\mathcal{D}_{\rm PL}^{tr_{1}},...,\mathcal{D}_{\rm PL}^{tr_{N}}\}$ $\mathcal{D}^{te}_{xy}$, $\{\mathcal{D}^{tr_{1}}_{xy},...,\mathcal{D}^{tr_{N}}_{xy}\}$
Supervised fine-tuning (SFT). SFT involves employing a supervised learning objective, specifically maximum likelihood estimation (MLE), to fine-tune a pretrained LLM on a task-specific dataset, denoted as $\mathcal{D}_{xy}$. The training objective is to obtain a language policy $\pi_{\theta}$ that satisfies
$\min_{\theta}-\mathbb{E}_{(x,y)\sim\mathcal{D}_{xy}}\left[\log\pi_{\theta}(y|x)\right]$ (1)
RLHF with reward modeling. The standard RLHF framework [37, 26] comprises two main phases: (i) learning an RM for a specific task, and (ii) fine-tuning the policy with the learned RM. During reward learning, given a preference distribution $\mathcal{D}_{\rm PL}$, an RM $r_{\phi}$ is trained to score preference of $y$ to be the answer of $x$, using a maximum likelihood estimation (MLE) loss function:
@@
where $\sigma$ denotes the logistic function. In the RL fine-tuning phase, given a prompt distribution $\mathcal{D}_{x}$, the optimization problem can be represented as follows,
$\max_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi_{\theta}(\cdot|x)||\pi_{\rm ref}(\cdot|x)\right)\right],$ (3)
where $\mathrm{D}_{\rm KL}(\cdot||\cdot)$ represents the Kullback-Leibler divergence, $\pi_{\rm ref}(\cdot|x)$ denotes a reference distribution, and $\beta$ is the balancing coefficient. Previous work [31, 30] uses an SFT model as the $\pi_{\rm ref}$ to impose constraints on the similarity between the predictions of the RLHF model and those of the initial SFT model. We define $\pi_{\rm ref}$ as the target distribution to address the policy drift challenge, as detailed in §3.
Problem Setting. We compare three PL methods in Tab. 1. To ensure a fair comparison, we consider a transfer learning scenario where all methods employ prompting from the test distribution $\mathcal{D}^{te}_{y|x}$ during testing. In contrast to supervised fine-tuning (SFT), preference learning (PL) utilizes additional preference data for training, as demonstrated in previous works [10, 26, 36, 22, 30]. In the ID PL setting, both RLHF-PPO [10, 26] and xDPO [30, 2] relay on the ID preference data $\mathcal{D}^{te}_{\rm PL}$ for reward learning or policy optimization. This study concentrates on an OOD scenario, where we utilize preference data $\{\mathcal{D}_{\rm PL}^{tr_{1}},...,\mathcal{D}_{\rm PL}^{tr_{N}}\}$ and RL fine-tuning data $\mathcal{D}^{te}_{xy}$, $\{\mathcal{D}^{tr_{1}}_{xy},...,\mathcal{D}^{tr_{N}}_{xy}\}$ from the training distributions for preference learning.
@@ 3 Meta-Learning for Generalizing Reward Modeling
Table 2: Linkes among different fields of bilevel optimization in this work.
Bilevel Programming ID Preference Learning OOD Preference Learning (Meta-Learning)
outer variables RM $\phi$ generalized RM $\phi$
outer objective PL objective $\mathcal{L}_{\rm PL}(\phi,\theta_{*},\mathcal{D}_{\rm PL})$ (Eq. (7)) PL objective ${\bf L}(\phi)$ (Eq. (11))
inner variables policy $\theta$ policies of each distribution $\{\theta^{T}\}_{T\in\mathcal{D}_{\mathcal{T}}}$
inner objective FT objective $\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D}_{xy})$ (Eq. (10)) FT objectives for each distribution $\{\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D}^{T}_{xy})\}_{T\in\mathcal{D}_{\mathcal{T}}}$ (Eq. (12))
Our meta-learning approach for OOD generalization is built upon bilevel programming, [16, 20, 19]. In this section, we first introduce a bilevel optimization framework for ID preference learning (PL) in §3.1. Based on this, we introduce a meta-learning approach for OOD PL in §3.2. Tab. 2 outlines the connections among bilevel programming, ID preference learning, and meta-learning for OOD PL. Additionally, we introduce a gradient-based algorithm for optimizing the meta-learning objective in §3.3 and analyze the theoretical convergence rate in §3.4.
3.1 Preference Learning via Bilevel Programming
Bilevel programming. We formularize preference learning as a bilevel optimization problem (see e.g., [12, 34]):
@@ 3.1 Preference Learning via Bilevel Programming
${\rm s.t.}\ \theta_{*}\in\mathop{\arg\min}_{\theta}\mathcal{L}_{\rm FT}({\phi},{\theta};\mathcal{D}_{xy}),\vspace{-0.5cm}$ (5)
@@
where $\mathcal{L}_{\rm PL}$ represents the outer objective, and for every reward function $r_{\phi}$, $\mathcal{L}_{\rm FT}$ represents the inner objective. Note that $\theta_{*}$ denotes the optimal parameters of a policy w.r.t. an RM $r_{\phi}$ for minimizing the inner RL objective.
@@
Outer objective. We desire that the fine-tuned policy ${\theta}_{*}$, with the help of the reward function $r_{\phi}$, can align with human preferences. In particular, given a pair of answers for a prompt $x$ with preference ranking, $\nu=(y,y^{\prime},x:y\succ y^{\prime}|x)\sim\mathcal{D}_{\rm PL}$, we use the Bradley-Terry (BT) model [8] to compute the comparing preference of $p(y\succ y^{\prime}|x)$ w.r.t. a fine-tuned policy ${\theta_{*}}$ as follows,
$\vspace{-0.5cm}p_{*}(y\succ y^{\prime}|x)=\frac{\pi_{\theta_{*}}(y|x)}{\pi_{\theta_{*}}(y|x)+\pi_{\theta_{*}}(y^{\prime}|x)}\vspace{-0.5cm}$ (6)
@@
Given a task-specific preference distribution $\mathcal{D}_{\rm PL}$, we apply the maximum likelihood estimation (MLE) to define the outer objective of preference alignment w.r.t. the policy ${\theta_{*}}$ as follows,
@@
$\begin{split}&\min_{\phi}\mathcal{L}_{\rm PL}(\phi,\theta_{*};\mathcal{D}_{\rm PL})=\min_{{\phi}}\mathbb{E}_{\nu\sim\mathcal{D}_{\rm PL}}\left[\ell_{\rm PL}({\phi},{\theta}_{*},\nu)\right]\\ &{\rm s.t.}\ \ell_{\rm PL}\left({\phi},{\theta}_{*},\nu\right)=-\log\sigma\left(\log\pi_{\theta_{*}}(y|x)-\log\pi_{\theta_{*}}(y^{\prime}|x)\right)\end{split}$ (7)
@@
Inner objective. The inner-loop optimization objective consist with the RL objective of Eq. (3), and especially the reference policy is defined as the task-specific distribution to avoid the policy drift problem. In particular, given a task-specific distribution $\mathcal{D}_{xy}$, the reference distribution is represented as the conditional task-specific distribution, i.e., $\pi_{\rm ref}(\cdot|x):=\mathcal{D}_{y|x}(\cdot)$ for any $x\sim\mathcal{D}_{x}$. Then, the RL fine-tuning objective can be represented as
@@
$\vspace{-0.5cm}\mathop{\arg\max}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi_{\theta}(\cdot|x)||\mathcal{D}_{y|x}(\cdot)\right)\right]\vspace{-0.5cm}$ (8)
@@
However, it is impossible to directly optimize the above objective, since the distribution $\mathcal{D}_{y|x}(\cdot)$ is unknown. Thus, we consider an approximate solution to this objective. The derivation largely follows Peters and Schaal [29] and Peng et al. [28]. The major observation is that the objective is strictly concave and thus we can use the Karush-Kuhn-Tucker (KKT) conditions [7] to obtain a globally optimal solution to Eq. (8), such that for any $x,y$,
$\vspace{-0.5cm}\pi_{*}(y|x)=\frac{1}{Z(x)}\mathcal{D}_{y|x}(y)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right),\vspace{-0.5cm}$ (9)
@@
where $Z(x)$ represents the partition function and $\beta$ represents the reward controlling factor. A detailed derivation is available in Appendix 0.A. We focus on a parameterized estimator (e.g., a neural network) to tackle a parameter estimation problem $\min_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}_{x}}[\mathrm{D}_{\rm KL}(\pi_{*}(\cdot|x)||\pi_{\theta}(\cdot|x))]$. Thereby, through sampling $z=(x,y)\sim\mathcal{D}_{xy}$, the inner objective can be represented by logical regression,
@@
$\vspace{-0.5cm}\begin{split}\mathop{\arg\min}_{\theta}\mathcal{L}_{\rm FT}(\phi,\theta,\mathcal{D}_{xy})&=\mathop{\arg\min}_{\theta}\mathbb{E}_{x\sim\mathcal{D}_{x}}\mathbb{E}_{y\sim\mathcal{D}_{y|x}}\left[\ell_{\rm FT}({\phi},{\theta},z)\right]\\ \qquad\qquad\qquad\qquad\quad&=\mathop{\arg\min}_{\theta}\mathbb{E}_{z\sim\mathcal{D}_{xy}}\left[\ell_{\rm FT}({\phi},{\theta},z)\right]\\ {\rm s.t.}\ \ell_{\rm FT}({\phi},{\theta},z)&=-\log\pi_{\theta}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\end{split}\vspace{-0.5cm}$ (10)
3.2 Meta-Learning for Out-of-Distribution Preference Learning
Figure 2: Overall training process. Meta-training optimizes the RM $\Phi$ using $K$ iterations of SGD. During testing, the RM $\Phi_{K}$ is utilized to optimize the test policy $\theta_{*}$.